OpenStudy (anonymous):

How would you solve: 3y-9≥2x

OpenStudy (anonymous):

Are you trying to get it into Slope Intercept Form?(Y = mx + b)?

OpenStudy (anonymous):

yeahh

zepdrix (zepdrix):

similar to the last one chloe? did you find that graph ok? :o

OpenStudy (anonymous):

yeah... It was okay. I just need more help. im a silly clueless girl.

OpenStudy (anonymous):

OK so treat the >= sign as just an = until the end of the problem. So... 3y-9 = 2x +9 +9 -------------- 3y = 2x +9 /3 /3 y = 2/3x + 3 BUT dont forget your sign.. Its really y >= 2/3x + 3 The only time you dont do this is when you are dividing or multiplying a negative, so be careful.

OpenStudy (anonymous):